. If one root of the quadratic equation ૛࢞૞ − ૛࢞ + ࢑ ૙ = is the reciprocal of the other

then find ࢑ .


1
Expert's answer
2021-09-01T11:41:07-0400
"ax^2+bx+c=0, a\not=0"

Given that one root of the quadratic equation is the reciprocal of the other


"x_1x_2=1"

Then


"D=b^2-4ac\geq0"

"x_1x_1=\dfrac{c}{a}=1=>c=a"

"b^2-4a^2\geq0"

"|b|\geq2|a|"



"a\not=0, |b|\geq2|a|, c=a."